Use generating functions to solve the recurrence relation an = 4an−1 − 4an−2 +n2

, where a0 = 2, a1 = 5.


1
Expert's answer
2021-12-14T01:18:12-0500
"a_n - 4a_{n\u22121} + 4a_{n\u22122} = n^2""\displaystyle\sum_{n=2}^{\infin}a_nx^n-4\displaystyle\sum_{n=2}^{\infin}a_{n-1}x^{n}+4\displaystyle\sum_{n=2}^{\infin}a_{n-2}x^{n}=\displaystyle\sum_{n=2}^nn^2x^n"

From the table


"\displaystyle\sum_{n=0}^nn^2x^n=0+x+2^2x^2+3^2x^3+...=\dfrac{x(x+1)}{(1-x)^3}""G(x)-2-5x-4x(G(x)-2)+4x^2G(x)""=\dfrac{x(x+1)}{(1-x)^3}-x""G(x)(1-4x+4x^2)=\dfrac{x(x+1)}{(1-x)^3}-4x+2""G(x)=\dfrac{x^2+x}{(1-x)^3(1-2x)^2}+\dfrac{2}{1-2x}""\dfrac{x^2+x}{(1-x)^3(1-2x)^2}=\dfrac{A}{(1-x)^3}+\dfrac{B}{(1-x)^2}+\dfrac{C}{1-x}""+\dfrac{D}{(1-2x)^2}+\dfrac{E}{1-2x}"





"A(1-2x)^2+B(1-x)(1-2x)^2""+C(1-x)^2(1-2x)^2+D(1-x)^3""+E(1-x)^3(1-2x)=\dfrac{x^2+x}{(1-x)^3(1-2x)^2}""x=0:A+B+C+D+E=0""x=-1:9A+18B+36C+8D+24E=0""x=1:A=2""x=1\/2:D=6""x=2:9A-9B+9C-D+3E=6"



"A=2, B=-3, C=-3, D=6, E=-2"




"G(x)=\dfrac{2}{(1-x)^3}-\dfrac{3}{(1-x)^2}-\dfrac{3}{1-x}""+\dfrac{6}{(1-2x)^2}-\dfrac{2}{1-2x}+\dfrac{2}{1-2x}""\def\arraystretch{1.5}
 \begin{array}{c:c}
 \dfrac{2}{(1-x)^3} & 2\dbinom{n+2}{2} \\ \\
 -\dfrac{3}{(1-x)^2} & -3(n+1)\\
 \\
 -\dfrac{3}{1-x} & -3 \\ \\
\dfrac{6}{(1-2x)^2} & 6(n+1)(2^n)
\end{array}""a_n=2\dbinom{n+2}{2}-3(n+1)-3+6(2^n)(n+1)"